OpenStudy (anonymous):

__ pi/3 sin pi/6=1/2(sin pi/2-sin pi/6) Im confused?

OpenStudy (callisto):

Perhaps this identity helps? \[cosAsinB = \frac{1}{2}[sin(A+B)-sin(A-B)]\]

OpenStudy (anonymous):

but i'm still confused on the question?

OpenStudy (callisto):

Now, you have two angle. Consider the left side, A = pi/3 B = pi/6 Try to apply this in the identity, and you'll see something